OpenStudy (anonymous):

find the equation of a line (2,-2) m=-5/3

OpenStudy (saifoo.khan):

y = mx + c -2 = 2(-5/3) + c c = -2 + 10/3 c = 4/3

OpenStudy (saifoo.khan):

y = -5x/3 + 4/3 3y = -5x + 4